Find three consecutive even integers such that the product of the first and the third is 24 less than 9 times the second.
Answer by mananth(16946)

let the consequitive even numbers be x,x+2,x+4
..
x(x+4)=9(x+2)-24
x^2+4x=9x+18-24
x^2+4x-9x-18+24=0
x^2-5x+6=0
x^2-3x-2x+6=0
x(x-3)-2(x-3)=0
(x-3)(x-2)=0
x=3 OR 2
2,4,6
...
m.ananth@hotmail.ca